OpenStudy (anonymous):

How can I find the slope after I find the derivative? For instance, f(x) = 2x^2 - 3x - 5, x = 0. I solved that question using the derivative formula and I got the answer 4x - 2. Now, how do I find the slope?

OpenStudy (amistre64):

the derivative is the equation of the slope at any given point

OpenStudy (amistre64):

y' = 4x - 2; when x=0, then y' = ??

OpenStudy (anonymous):

It's confusing me, because it says the answer is m= -3

OpenStudy (amistre64):

i see, you dint derive it correctly, check it out again

OpenStudy (amistre64):

2x^2 - 3x - 5 ; what do we derive it to?

OpenStudy (anonymous):

Compute the derivative of the given function and find the slope of the line that is tangent to its graph for the specified value of the independent variable. f(x) = 2x^2-3x-5; x =0

OpenStudy (amistre64):

and your derivative is?

OpenStudy (anonymous):

Thats the question. I used the derivative formula and I got the answer 4x -2

OpenStudy (amistre64):

-3x does not derive down to -2

OpenStudy (anonymous):

wow, thank you! You just made me realize what I wasn't seeing. I see it for all the problems now. So once I get the answer. The slope is the derivative when I plug in the X

OpenStudy (amistre64):

correct: y' = 4x - 3 , at x=0; y' = slope = -3

OpenStudy (anonymous):

Are you able to help me out in one more thing?

OpenStudy (amistre64):

maybe, depends on if i know it or not :)

OpenStudy (anonymous):

I need to find the tangent

OpenStudy (amistre64):

the tangent, or the equation of the tangent line?

OpenStudy (anonymous):

I have to compute the derivative of the given function then find the equation of the line that is tangent to its graph for the specified value x = c

OpenStudy (amistre64):

do you recall the point slope form of a line? from algebra?

OpenStudy (anonymous):

m = y - y1 = m(x - x1)

OpenStudy (anonymous):

sorry I didnt men to put M at the front

OpenStudy (amistre64):

yes, but not the first part: y - yo = m(x-xo) lets work on this a little y - yo = mx - mxo y = mx -mxo + yo we know m = y' = -3 when x=0 so lets fill those in y = -3x +3(9) + yo we get yo form the original equation: y = 4x^2 -3x -5 ; at x=0

OpenStudy (amistre64):

y0 = -5 in that case so lets finish it y = -3x +3(0) - 5 the equation of the tangent line at x=0 is then y = -3x -5

OpenStudy (amistre64):

its either that, or they want the point slope form itself, depends on the material i spose

OpenStudy (anonymous):

What's y0? a y then zero?

OpenStudy (amistre64):

its a slip of the typing :) the o and 0 are so close on the keyboard ...

OpenStudy (amistre64):

that and i also mistyped y = -3x +3(9) + yo ^^^ shoulda been 3(0) , but i hit the wrong key again

OpenStudy (anonymous):

Ok, I get it. THANK YOU :)

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

HOLD ON, M is the slope and y?

OpenStudy (amistre64):

m is traditionally slope, but m = y' so dont get hung up on the letters

OpenStudy (amistre64):

the important thing is to recognize what it is your after, and then adapt your own method of getting there

OpenStudy (anonymous):

Oh so mx is just the X they give me

OpenStudy (anonymous):

It says compute the derivative of the given function and then find the equation of the line that is tangent to its graph for the specified value x=c. f(x) = 2; c=13 another problem is: f(x) 7-2x; c=5

OpenStudy (amistre64):

lets try to organise it this way: \[y_o=2x_o^2-3x_o-5;\ at\ x_o=0,\ y_o=-5\]\[y_o' = 4x_o-3;\ at\ x_o=0,\ y_o'=-3\] now form the tangent equation \[y-y_o=y_o'(0)(x-x_o)\] \[y+5=-3(x-0)\]and arrange as needed

OpenStudy (anonymous):

is that from the problem f(x) = 2x^2-3x-5; x=0 ?